Hey the correct answer is given for both, but can you please do the steps on how to get to the answer, Im confused on how to solve? Thank you!

3. At least what sample size is necessary in order to obtain a 90% confidence interval of length 6, when SD = 10? A. 8 B. 16 C. 60 D. 15 E. 61 Answer: E

4. With which combination of confidence level and sample size will the corresponding confidence interval be largest? A. 90%, n = 25 B. 99%, n = 50 C. 80%, n = 20 D. 95%, n = 25 E. 98%, n = 50 Answer: D
Answer #1

Confidence interval in given by (x + E) (X - E) where, E= Margin of error. length of interval = (7+E) - (7-E)= 6 2E=6 E=6 9